E . The red light of wavelength 750 nm and blue light of 450 nm are illuminating a film of n=1.2 surrounded by the air.

a). What is the minimum thickness of this film which will provide the maximum reflection for both these wavelengths? Give answer in nm. (IT IS A SINGLE ANSWER, NOT TWO INDIVIDUAL)

b). The next after minimum thickness of the film which provides maximum reflection for both wavelengths in the previous question will be? Give answer in nm. IT IS A SINGLE ANSWER, NOT TWO INDIVIDUAL)

c). The minimum thickness of a film which provides the total transmission for wavelengths from question E? Give answer in nm. (IT IS A SINGLE ANSWER, NOT TWO INDIVIDUAL)

d). The next after minimum thickness which provides maximum transmission for both wavelengths from question E ? Give answer in nm. (IT IS A SINGLE ANSWER, NOT TWO INDIVIDUAL)
Answer #1

M2 = 1.2 (film) In case of reflection (nean, ) I (air) for maximum reflection! tata waredagth inside film 27-(m) as n film finfilm analm 4x1.2 150x100m uxl.2 t. 122*9 (i) at = (m-1) for mal 2t=1 ta j t = (a) (for red light) t = 156.25x10 m = 0.156816So, thickness required in It =0.4683m) (it) for total focansmistion:- om=1, 2, 3 . [2t = min 2 tnfilm -md (a) for mal tad ta